OpenStudy (anonymous):

A cup of yogurt cost $2.60 plus ten cents per topping. A cup of ice cream costs $2.00 plus twenty five cents per topping. How many toppings are needed so that the two desserts cost the same?

OpenStudy (anonymous):

2.6+0.1x = 2+0.25x x=4 4 toppings are needed